# xTrimoPGLM: Unified 100B-Scale Pre-trained Transformer for Deciphering the Language of Protein

### v1

@@
[…]
${}^{1}$ioMap Research ${}^{2}$singhua University ${}^{3}$
Equal Contribution. Emails: cb21@mails.tsinghua.edu.cn, xingyi@biomap.comWork done during interning at BioMap Research, California, USA.The project leader at BioMap Research, California, USA.The corresponding authors. Emails: songle@biomap.com, jietang@tsinghua.edu.cn
Protein language models have shown remarkable success in learning biological information from protein sequences. However, most existing models are limited by either autoencoding or autoregressive pre-training objectives, which makes them struggle to handle protein understanding and generation tasks concurrently. We propose a unified protein language model, xTrimoPGLM, to address these two types of tasks simultaneously through an innovative pre-training framework. Our key technical contribution is an exploration of the compatibility and the potential for joint optimization of the two types of objectives, which has led to a strategy for training xTrimoPGLM at an unprecedented scale of 100 billion parameters and 1 trillion training tokens. Our extensive experiments reveal that 1) xTrimoPGLM significantly outperforms other advanced baselines in 18 protein understanding benchmarks across four categories. The model also facilitates an atomic-resolution view of protein structures, leading to an advanced 3D structural prediction model that surpasses existing language model-based tools. 2) xTrimoPGLM not only can generate de novo protein sequences following the principles of natural ones, but also can perform programmable generation after supervised fine-tuning (SFT) on curated sequences. These results highlight the substantial capability and versatility of xTrimoPGLM in understanding and generating protein sequences, contributing to the evolving landscape of foundation models in protein science.
@@ Introduction
Proteins play vital roles in the sustenance, growth, and defense mechanisms of living organisms. They provide structural support for many essential biological processes such as synthesizing enzymes, facilitating transportation, regulating gene expression, and contributing to immune function. Therefore, understanding the biological information encoded within proteins is crucial for unraveling the intricate workings of life and advancing fields such as medicine and biotechnology (?, ?). As protein sequences serve as the blueprint for protein structure and function (?), pre-trained techniques on sequences, known as Protein Language Models (PLMs), e.g., the family of ESM models (?, ?), ProtTrans (?), ProGen (?), etc., offer a powerful tool for characterizing the properties and distributions of general protein sequences. These models are trained on large-scale protein datasets (?, ?, ?, ?) that encompass billions of sequences, allowing them to capture evolutionary patterns and sequence features that are inherent in protein structures. As a result, these models achieve state-of-the-art results in predicting protein functions and structures (?, ?) or generating novel sequences with faithful three-dimensional structures (?, ?).
@@
It is worth noting that different categories of protein-related tasks necessitate divergent outputs from PLMs, such as protein understanding tasks call for PLMs to yield accurate residue-level or protein-level representations, while protein design tasks depend heavily on the potent generation capabilities of PLMs. Despite these varying outputs, all tasks reveal a consistent underlying dependency among protein sequences (?, ?, ?), which suggests the possibility of characterizing these tasks within one unified framework, potentially mitigating the disparity between task types and further augmenting the modeling power of PLMs. Unfortunately, existing PLMs are designed to address specific tasks depending on their pre-training framework. This presents a significant challenge to selecting appropriate PLMs for specific task types. Consequently, we explore the feasibility of integrating tasks of understanding and generation, dictated by autoencoding and autoregressive pre-training objectives, respectively, into one unified framework. This unified approach aims to encapsulate the intricate dependencies inherent in protein sequences, potentially resulting in more versatile and robust protein foundation models.
@@
Large Language Models (LLMs) have explored the revenue of developing unified pre-training paradigms. However, these studies typically adopt analogous training patterns. For instance, all pre-training objectives are commonly optimized using either the BERT-style (?, ?) or GPT-style regime (?). A balanced approach incorporating both bi-directional auto-encoding and uni-directional auto-regressive objectives could fulfill the requirements of unified PLMs, yet the feasibility of such integration remains an open question. Practically, the current landscape of natural language processing is dominated by generative models, which afford various types of tasks via mapping task labels into a unified text space for zero/few-shot learning (?) or instruction-tuning (?, ?). However, this capability is currently beyond the reach of PLMs. In practice, applications of protein modeling still rely on the bridging of representations with downstream task-specific labels, such as discrete values of categories or continuous values of 3D coordinates (?, ?). These tasks heavily rely on bi-directional auto-encoding training to tackle protein understanding tasks. Consequently, this highlights the need for a unified model that incorporates both training objectives.
Key Results
@@
Our first key result is the xTrimo Protein General Language Model (xTrimoPGLM), a unified pre-training framework and foundation model designed for various protein-related tasks, including understanding and generation (or design). The model differs from previous encoder-only (e.g., ESM) or causal decode-only (e.g., ProGen) protein language models by leveraging the General Language Model (GLM) (?) as the backbone for its bidirectional attention and auto-regressive objective. To enhance the representation capacity of xTrimoPGLM, we further introduce the Masked Language Model (MLM) objective to the bidirectional prefix region, building upon the generation ability encapsulated within the GLM objective. Additionally, we compiled a large pre-training dataset, comprising approximately 940 million unique protein sequences with roughly 200 billion residues, and trained a model with 100 billion parameters over 1 trillion tokens over a cluster of 96 NVIDIA DGX machines each with 8$\times$×A100 GPU cards. To further justify the superiority of the xTrimoPGLM framework, we train an antibody-specific model (xTrimoPGLM-Ab) using the same framework with 1 billion parameters. This model sets a new record in predicting antibody naturalness and structures, both essential to the field of antibody-based drug design, and demonstrates a significantly faster inference speed than AlphaFold2, see supplementary material (SM) section 2.
@@
The second key result is a significant enhancement in the realm of protein understanding with the xTrimoPGLM-100B as a protein foundation model. By conducting extensive empirical experiments with linear probing and advanced fine-tuning techniques, we elevated the performance benchmarks in this domain. xTrimoPGLM-100B has significantly surpassed previous state-of-the-art (SOTA) methods in 15 out of 18 tasks, covering a comprehensive range of areas including protein structure, interactions, functionality, and developability (Figure 2A). We also illustrate that xTrimoPGLM achieves lower Perplexity (PPL) on two Out-Of-Distribution (OOD) protein sets over other models (Figure 1B). These results empirically validate the scaling behavior, demonstrating that larger models commonly tend to yield better performance (Figure 1C & Figure 2B).
@@
The third key result is the development of a high-performance 3D structural prediction tool based on xTrimoPGLM. Inspired by methodologies similar to those in ESMFold, merge folding modules with a protein language model, thereby refining protein structure training. Our version named xTrimoPGLM-Fold (xT-Fold for short) has shown promising results with impressive TM-scores in both CAMEO (n=194) and CASP15 (n=56) protein benchmarks. Additionally, we optimized xT-Fold through 4-bit quantization, enhancing its performance and efficiency, which makes xT-Fold a leading option in PLM-based structure prediction tools. As a result, xT-Fold achieves a 5-point TM-score increase over ESMFold in the CASP15 dataset, coupled with a faster inference speed across various scenarios (Figure 3).
@@
The final result demonstrates that xTrimoPGLM showcases an extraordinary ability to generate de novo protein sequences. These sequences not only exhibit diverse structures closely akin to natural counterparts, as evidenced by a median sequence identity of just 11.7% (Figure 4), but can also be tailored towards specific structural and biophysical properties through supervised fine-tuning (Figure 5 & Figure 6). This “super alignment” capability of xTrimoPGLM underscores its potential as a programmable model for exploring and synthesizing the vast protein space.
@@
Figure 1: Comprehensive Insights into xTrimoPGLM. A. xTrimoPGLM Development Phases: illustrates the pre-training and fine-tuning stages of xTrimoPGLM, combining GPT-style and BERT-like objectives, and applying learned embeddings in protein understanding and design tasks. B. Perplexity Comparison: xTrimoPGLM shows lower perplexity than leading models like ESM2 and ProGen2-xlarge in evaluations on two distinct out-of-distribution datasets, indicating its advanced performance. C. Scaling Law Analysis: examines the scalability of models from 1 million to 1 billion parameters, trained with 100 billion tokens, demonstrating xTrimoPGLM-100B’s efficiency through a power law fit of training losses against computational resources.
@@
Quantification Scaling Law of xTrimoPGLM. The scaling law is a crucial concept for understanding the performance of large language models (LLMs) during pre-training. It suggests a power-law relationship between a model’s performance, typically measured by cross-entropy test loss $L(C)$, adheres to a power-law relationship with the computer resources $C$ used in training. To investigate the scaling law for xTrimoPGLM in the context of protein data, we employ the formula $L=a\times C^{b}$ where $C$ as outlined by (?, ?). We quantify $C$ in terms of training floating point operations (FLOPs), using PF-days (PetaFLOP-days) as the unit of measure, where one PF-day$=$ 8.64 × $10^{19}$ FLOPs. We extract a range of $(C,L(c))$ data points from the loss trajectories of models varying from 1 million to 1 billion parameters, each trained with 100 billion tokens. The resultant scaling curve, fitted to these data points, is depicted in Figure 1C. Remarkably, this curve demonstrates a close alignment with the actual training loss observed for the 100 billion parameter model of xTrimoPGLM, trained with the same volume of tokens. This observation substantiates the model’s adherence to the anticipated scaling law, thereby reinforcing the credibility of this principle as a pivotal guideline in the development of large-scale protein language models.
@@
xTrimoPGLM Achieves Low Perplexity on Out-Of-Distribution Protein Sequences. Perplexity is a metric in language modeling that quantifies the uncertainty of a probability model in predicting a text sample, where lower values signify greater predictive accuracy. For a comprehensive evaluation, we construct two OOD datasets using a specific process: (1) Sampling an extensive collection of protein sequences from UniProt that were uploaded after January 2023, which is also the cutoff date for the training data. And (2) Filtering these sequences based on sequence identity thresholds of 0.5 and 0.9 with the pre-training datasets. Each OOD dataset comprises approximately 10,000 protein sequences. The results are demonstrated in Figure 1B, xTrimoPGLM-100B achieves perplexity scores of 10.81 and 6.70 on the 0.5 and 0.9 sequence identity datasets, respectively, outperforming 15 billion parameters ESM2-15B (10.98 and 7.25). Similarly, against ProGen2-xlarge with 6.4 billion parameters, xTrimoPGLM recordes scores of 13.35 and 8.78, compared to ProGen2-xlarge’s 14.30 and 9.75. We observed that large-scale models are more sample efficient, reaching these perplexity levels with substantially less training data: less than 500 million tokens to match ESM2 trained on 1 trillion tokens, and less than 150 billion tokens to parallel ProGen2-xlarge trained with 350 billion tokens, even if the learning rate schedule has not yet ended, resulting in an overestimated loss value at this stage.
Extensive Benchmarks of 18 Protein Understanding Tasks
@@
To comprehensively assess the understanding capabilities of xTrimoPGLM-100B, we benchmark it against 18 downstream protein-related tasks. These tasks span four primary categories: protein structure, developability, interactions, and functions, as detailed in Figure 2A. A thorough description of these datasets can be found in SM Section 1.8. We compare xTrimoPGLM-100B with two other protein language models in the field, ESM2-150M and ESM2-15B (?), to provide a well-rounded evaluation. Our assessment methodology encompasses two distinct approaches to evaluate the effectiveness of the models’ representations:
@@
  • Fine-tuning with LoRA. Considering the constraints of GPU memory, full-scale fine-tuning is not feasible for models with parameters on the scale of 100 billion. Hence, we employ Low-Rank Adaptation (LoRA) (?) as a parameter-efficient alternative. This technique involves freezing the pre-trained model’s weights and adding trainable low-rank matrices to each transformer layer. LoRA significantly reduces the number of trainable parameters for downstream tasks while preserving the adaptability of the learned representations. The fine-tuning architecture and settings are similar to those in MLP probing, with only the transformer’s ${W_{q},W_{k},W_{v},W_{o}}$ parameters being fine-tuned.
Figure 2: The Performance of Protein Understanding Benchmark. A. Visualization of Model Performance across all Tasks. B. The scaling trend between the computational complexity of model training, quantified by PF-days, and the model performance across four distinctive task categories, each encompassing 3-4 subtasks. Each data point symbolizes the mean performance metric for a specific task category, where Pb stands for Probing and Ft for Fine-tuning with LoRA techniques. The labels E150M, E650M, E3B, E15B, and xT100B represent ESM2-150M, ESM2-650M, ESM2-3B, ESM2-15B, and xTrimoPGLM-100B, respectively. C. Correlations between the Pre-Training validation loss measured by MLM objective and the downstream task performance. To facilitate comparison, we normalize this performance by subtracting the mean value and dividing by the standard deviation. Most tasks (44% Gradually Increasing, 28% Grokking, and 16% Not Strictly Linear) obey the trends that the task performance demonstrates the positive correlations with the ongoing pre-training process.
@@
Figure 2A provides a comprehensive visualization of performance across all benchmarked protein-related tasks. Distinct colors and shapes in the figure represent different evaluation strategies and models, respectively. The ESM2-150M model, being relatively smaller, serves as an indicator to assess the difficulty of these tasks. The performance distribution highlights the inherent relationships between the complexity of tasks and the advantages brought by the scale of the model. The distribution of performance across tasks underscores the relationship between task complexity and the benefits derived from the scale of the model. In more complex tasks, such as Contact Map prediction (Protein Structure category), Fluorescence (Protein Function), Metal Binding (Protein Interaction), and Stability (Protein Development), the larger models (xTrimoPGLM-100B and ESM2-15B) significantly outperform the smaller ESM2-150M. This disparity in performance highlights the necessity for more advanced models to effectively address complex tasks. Conversely, for simpler tasks, like Antibiotic Resistance (Protein Function category), the performance gap between the large and small models is notably smaller. This observed pattern suggests that larger models are more adept at capturing the intrinsic evolutionary information of protein sequences. Consequently, as models scale up, they exhibit marked improvements in performance, particularly for complex tasks. The application of LoRA consistently enhances overall task performance compared to the static Probing method, which limits the capacity of pre-trained models (as illustrated in Figure2B). LoRA’s efficiency lies in its ability to refine and utilize key features without significantly increasing the number of trainable parameters (less than 1% increase with LoRA rank set to 8). The empirical findings also highlight a scaling trend in task performance with supervised fine-tuning, suggesting a strong correlation between model scale and performance. The following sections will delve deeper into the varying scaling behaviors observed across different types of tasks.
@@
Scaling Behaviors of Downstream Tasks. We extend the investigation of scaling laws to downstream tasks. We examine the correlation between computational complexity (measured by PF-days) during model training and performance across four distinct categories of protein-related tasks. Each category encompasses 3-4 subtasks, analyzed using both Linear Probing with MLP (Pb) and Fine-tuning with LoRA (Ft) techniques (Figure 2B). While the intensity of the scaling effect differs among task types, a common trend observed from pre-training extends to these downstream tasks: an exponential increase in computational resources during pre-training correlates with linear improvements in task performance. This extension of the scaling law to downstream tasks in protein modeling is a novel observation.
@@
A prevalent assumption posits that large language models, due to their higher computational complexity, universally outperform smaller models across all tasks. To validate this, we take a deep insight into the correlations between the downstream task performance with the ongoing pre-training process (measured by the MLM validation loss), as shown in Figure 2C. We observe that most tasks demonstrate positive correlations. Specifically, 44% show a gradual increase in performance, 28% exhibit a “Grokking” phenomenon, and 16% do not follow a strictly linear pattern. The ’Grokking’ phenomenon mirrors emergent abilities seen in large NLP models (?). It occurs when models initially overfit the training data, especially when task datasets have limited overlap with pre-training data. As the model’s knowledge base expands, it begins to apply its understanding to out-of-distribution (OOD) scenarios, resulting in a sudden increase in test data performance. Thus the performance of the test data suddenly increased, denoted as the Grokking phenomenon. However, 16% of tasks indicate a potential negative impact from increased computational resources, suggesting an inverse scaling effect.

### v2

@@
Bo Chen2  , Xingyi Cheng [^1] footnotemark: 1
Pan Li1, Yangli-ao Geng [^2] footnotemark: 2, Jing Gong1, Shen Li1, Zhilei Bei [^2] footnotemark: 2, Xu Tan1,
Boyan Wang [^2] footnotemark: 2, Xin Zeng1, Chiming Liu1, Aohan Zeng2, Yuxiao Dong2
Jie Tang244footnotemark: 4, Le Song1,3
1BioMap Research 2Tsinghua University 3MBZUAI
Equal Contribution. Emails: cb21@mails.tsinghua.edu.cn, derrickzy@gmail.comWork done during interning at BioMap Research, California, USA.The project leader at BioMap Research, California, USA.The corresponding authors. Emails: dasongle@gmail.com, jietang@tsinghua.edu.cn
Protein language models have shown remarkable success in learning biological information from protein sequences. However, most existing models are limited by either autoencoding or autoregressive pre-training objectives, which makes them struggle to handle protein understanding and generation tasks concurrently. We propose a unified protein language model, xTrimoPGLM, to address these two types of tasks simultaneously through an innovative pre-training framework. Our key technical contribution is an exploration of the compatibility and the potential for joint optimization of the two types of objectives, which has led to a strategy for training xTrimoPGLM at an unprecedented scale of 100 billion parameters and 1 trillion training tokens. Our extensive experiments reveal that 1) xTrimoPGLM significantly outperforms other advanced baselines in 18 protein understanding benchmarks across four categories. The model also facilitates an atomic-resolution view of protein structures, leading to an advanced 3D structural prediction model that surpasses existing language model-based tools. 2) xTrimoPGLM not only can generate de novo protein sequences following the principles of natural ones, but also can perform programmable generation after supervised fine-tuning (SFT) on curated sequences. These results highlight the substantial capability and versatility of xTrimoPGLM in understanding and generating protein sequences, contributing to the evolving landscape of foundation models in protein science. Trained weight for the xTrimoPGLM model, and downstream datasets are available at https://huggingface.co/proteinglm.
@@ Introduction
Proteins play vital roles in the sustenance, growth, and defense mechanisms of living organisms. They provide structural support for many essential biological processes such as synthesizing enzymes, facilitating transportation, regulating gene expression, and contributing to immune function. Therefore, understanding the biological information encoded within proteins is crucial for unraveling the intricate workings of life and advancing fields such as medicine and biotechnology  (?, ?, ?). As protein sequences serve as the blueprint for protein structure and function  (?), pre-trained techniques on sequences, known as Protein Language Models (PLMs), e.g., the family of ESM models  (?, ?), ProtTrans  (?), PROGEN  (?), etc., offer a powerful tool for characterizing the properties and distributions of general protein sequences. These models are trained on large-scale protein datasets  (?, ?, ?) that encompass billions of sequences, allowing them to capture evolutionary patterns and sequence features that are inherent in protein structures. As a result, these models achieve state-of-the-art results in predicting protein functions and structures  (?, ?, ?) or generating novel sequences with faithful three-dimensional structures  (?, ?).
@@
It is worth noting that different categories of protein-related tasks necessitate divergent outputs from PLMs, such as protein understanding tasks call for PLMs to yield accurate residue-level or protein-level representations, while protein design tasks depend heavily on the potent generation capabilities of PLMs. Despite these varying outputs, all tasks reveal a consistent underlying dependency among protein sequences  (?, ?), which suggests the possibility of characterizing these tasks within one unified framework, potentially mitigating the disparity between task types and further augmenting the modeling power of PLMs. Unfortunately, existing PLMs are designed to address specific tasks depending on their pre-training framework. This presents a significant challenge to selecting appropriate PLMs for specific task types. Consequently, we explore the feasibility of integrating tasks of understanding and generation, dictated by autoencoding and autoregressive pre-training objectives, respectively, into one unified framework. This unified approach aims to encapsulate the intricate dependencies inherent in protein sequences, potentially resulting in more versatile and robust protein foundation models.
@@
Large Language Models (LLMs) have explored the revenue of developing unified pre-training paradigms. However, these studies typically adopt analogous training patterns. For instance, all pre-training objectives are commonly optimized using either the BERT-style  (?) or GPT-style regime  (?). A balanced approach incorporating both bi-directional auto-encoding and uni-directional auto-regressive objectives could fulfill the requirements of unified PLMs, yet the feasibility of such integration remains an open question. Practically, the current landscape of natural language processing is dominated by generative models, which afford various types of tasks via mapping task labels into a unified text space for zero/few-shot learning  (?) or instruction-tuning  (?, ?). However, this capability is currently beyond the reach of PLMs. In practice, applications of protein modeling still rely on the bridging of representations with downstream task-specific labels, such as discrete values of categories or continuous values of 3D coordinates  (?, ?). These tasks heavily rely on bi-directional auto-encoding training to tackle protein understanding tasks. Consequently, this highlights the need for a unified model that incorporates both training objectives.
@@
In this work, we develop the first, to our knowledge, the xTrimo Protein General Language Model (xTrimoPGLM), a unified pre-training framework and foundation model that scales up to 100 billion parameters, designed for various protein-related tasks, including understanding and generation (or design). The model differs from previous encoder-only (e.g., ESM) or causal decode-only (e.g., PROGEN) protein language models by leveraging the General Language Model (GLM)  (?) as the backbone for its bidirectional attention and auto-regressive objective. To enhance the representation capacity of xTrimoPGLM, we further introduce the Masked Language Model (MLM) objective to the bidirectional prefix region, building upon the generation ability encapsulated within the GLM objective. Additionally, we compiled a large pre-training dataset, comprising approximately 940 million unique protein sequences with roughly 200 billion residues, and trained a model with 100 billion parameters over 1 trillion tokens over a cluster of 96 NVIDIA DGX machines each with 8$\times$×A100 GPU cards.
@@
xTrimoPGLM-100B demonstrates the significant enhancement in the realm of protein understanding. By conducting extensive empirical experiments with linear probing and advanced fine-tuning techniques, we elevated the performance benchmarks in this domain. xTrimoPGLM-100B has significantly surpassed previous state-of-the-art (SOTA) methods in 15 out of 18 tasks, covering a comprehensive range of areas including protein structure, interactions, functionality, and developability (Figure 2A). We also illustrate that xTrimoPGLM achieves lower Perplexity (PPL) on two Out-Of-Distribution (OOD) protein sets over other models (Figure 1B). These results empirically validate the scaling behavior, demonstrating that larger models commonly tend to yield better performance (Figure 1C and Figure 2B).
@@
xTrimoPGLM can serve as the base for developing a high-performance 3D structural prediction tool. Inspired by methodologies similar to those in ESMFold, merge folding modules with a protein language model, thereby refining protein structure training. Our version named xTrimoPGLM-Fold (xT-Fold for short) has shown promising results with impressive TM-scores in both CAMEO (n=194) and CASP15 (n=56) protein benchmarks. Additionally, we optimized xT-Fold through 4-bit quantization, enhancing its performance and efficiency, which makes xT-Fold a leading option in PLM-based structure prediction tools. As a result, xT-Fold achieves a 5-point TM-score increase over ESMFold in the CASP15 dataset, coupled with a faster inference speed across various scenarios (Figure 3).
@@
xTrimoPGLM also showcases an extraordinary ability to generate de novo protein sequences. These sequences not only exhibit diverse structures closely akin to natural counterparts, as evidenced by a median sequence identity of just 11.7% (Figure 4), but can also be tailored towards specific structural and biophysical properties through supervised fine-tuning (Figure 5 and Figure 6). This “super alignment” capability of xTrimoPGLM underscores its potential as a programmable model for exploring and synthesizing the vast protein space.
@@ Results
Figure 1: Comprehensive Insights into xTrimoPGLM. A. The pre-training and fine-tuning stages of xTrimoPGLM, combining BERT-style (blue and purple, for masking and predicting tokens) and GPT-style (green, from [S] to [E], for autoregressive generation) objectives. The prefix’s bidirectional attention facilitates protein understanding tasks like structure prediction, while the suffix supports both de novo and conditional protein design through sequence generation. B. xTrimoPGLM shows lower perplexity than other leading PLMs like ESM2 and PROGEN2-xlarge in evaluations on two distinct out-of-distribution datasets, indicating its advanced performance. C. The scaling behavior of xTrimoPGLM-series from 1 million to 1 billion parameters, trained with 100 billion tokens, demonstrating xTrimoPGLM-100B’s efficiency through a power law fit of training losses against computational resources.
@@
Quantification Scaling Law of xTrimoPGLM Family Models. The scaling law is a crucial concept for understanding the performance of large language models (LLMs) during pre-training. It suggests a power-law relationship between a model’s performance, typically measured by cross-entropy test loss $L(C)$, adheres to a power-law relationship with the computer resources $C$ used in training. To investigate the scaling law for xTrimoPGLM in the context of protein data, we employ the formula $L=a\times C^{b}$ where $C$ as approximated by $6ND$, where $N$ is the model size and $D$ is the pre-trained dataset size (set to 100 billion tokens in this case for all the xTrimoPGLM family models)   (?, ?). Thus the scaling behaviour between ($L(C)$) and total compute $C$ can also be viewed as the rule between ($L(C)$) and model scale ($N$). We quantify $C$ in terms of training floating point operations (FLOPs), using PF-days (PetaFLOP-days) as the unit of measure, where one PF-day$=$ 8.64 × $10^{19}$ FLOPs. We extract a range of $(C,L(C))$ data points from the loss trajectories of models varying from 1 million to 1 billion parameters, each trained with 100 billion tokens (Figure 1C). Remarkably, this curve demonstrates close alignment with the actual training loss observed for the 100 billion parameter model of xTrimoPGLM, trained with the same volume of tokens. This observation substantiates the model’s adherence to the anticipated scaling law, thereby reinforcing the credibility of this principle as a pivotal guideline in the development of large-scale protein language models.
@@
xTrimoPGLM Achieves Low Perplexity on Out-Of-Distribution Protein Sequences. Perplexity is a metric in language modeling that quantifies the uncertainty of a probability model in predicting a text sample, where lower values signify greater predictive accuracy. For a comprehensive evaluation, we construct two OOD datasets using a specific process: (1) Sampling an extensive collection of protein sequences from UniProt that were uploaded after January 2023, which is also the cutoff date for the training data. And (2) Filtering these sequences based on sequence identity thresholds of 0.5 and 0.9 with the pre-training datasets. Each OOD dataset comprises approximately 10,000 protein sequences (Figure 1B). xTrimoPGLM-100B achieves perplexity scores of 10.81 and 6.70 on the 0.5 and 0.9 sequence identity datasets, respectively, outperforming 15 billion parameters ESM2-15B (10.98 and 7.25). Similarly, against PROGEN2-xlarge with 6.4 billion parameters, xTrimoPGLM recorders scores of 13.35 and 8.78, compared to PROGEN2-xlarge’s 14.30 and 9.75. We observe that large-scale models are more sample efficient, reaching these perplexity levels with substantially less training data with respect to the corresponding pre-training objective: less than the actual 480 billion trained tokens for MLM pre-training (400 billion tokens at phases-1 and 80 out of 100 billion at phases-2) to match ESM2 trained on 1 trillion tokens, and less than actual 120 billion tokens for the GLM pre-training (120 out of 150 billion tokens at phase-2) to parallel PROGEN2-xlarge trained with 350 billion tokens, even if the learning rate schedule has not yet ended, resulting in an overestimated loss value at this stage. Noted that, the empirical studies in the Supplementary Section Empirical Analysis of Unified Training confirm that the GLM model, when continued from MLM pre-training, converges faster than GLM model pre-trained from scratch. Thus the initial 400 billion MLM tokens serve as a warm-up stage, enhancing the model’s ability to understand sequence distributions. This foundational understanding accelerates the convergence during the subsequent GLM pre-training phase to enhance the model’s generation capacity.
Figure 2: The Performance of Protein Understanding Benchmark. A. For the classification task, four metrics are employed (Supplementary Section 1) including TopL/5 accuracy (Contact map), accuracy (Fold classification, Secondary structure, Antibiotic resistance, Solubility, Localization, Metal ion binding), AUC (Peptide-HLA/MHC affinity, TCR-pMHC affinity, Clone CLF, Material production) and Matthews Correlation Coef. (Temperature stability). For the regression task, two metrics are used including the Spearman Correlation Coef. (Fluorescence, Fitness, Stability, Optimal temperature, Optimal PH) and the Pearson Correlation Coef. (Enzyme catalytic efficiency). B. The scaling trend between the computational cost of model training, quantified by PF-days, where one PF-day$=$ 8.64 × $10^{19}$ FLOPs, and the model performance Each data point symbolizes the mean performance metric for a specific task category (Pb for Probing and Ft for Fine-tuning with LoRA). E150M/650M/3B/15B, and xT100B represent ESM2-150M/650M/3B/15B, and xTrimoPGLM-100B, respectively. C. Correlations between the pre-training validation loss measured by MLM objective and the performance of the downstream tasks. To facilitate comparison, we normalize this performance by subtracting the mean value and dividing it by the standard deviation.
Evaluation on Protein Understanding benchmarks
@@
To comprehensively assess the understanding capabilities of xTrimoPGLM-100B, we benchmark it against 18 downstream protein-related tasks. These tasks span four primary categories: protein structure, developability, interactions, and functions (Figure 2A and Supplementary Section 1). Protein Structure: This category involves modeling the protein structure from the primary sequence, including secondary and tertiary structures. Protein Developability: This category focuses on the engineering characteristics of proteins, such as their stability and manufacturability. Protein Interaction: This category includes scenarios where proteins interact with other molecules, such as short peptides or other proteins. Protein Function: This category encompasses tasks that predict the intrinsic cellular features and activities of proteins, including enzyme catalytic efficiency. We compare xTrimoPGLM-100B with two other protein language models in the field, ESM2-150M and ESM2-15B  (?), to provide a well-rounded evaluation. Our assessment methodology encompasses two distinct approaches to evaluate the effectiveness of the models’ representations:
@@
  • Fine-tuning with LoRA. Considering the constraints of GPU memory, full-scale fine-tuning is not feasible for models with parameters on the scale of 100 billion. Hence, we employ Low-Rank Adaptation (LoRA)  (?) as a parameter-efficient alternative. This technique involves freezing the pre-trained model’s weights and adding trainable low-rank matrices to each transformer layer. LoRA significantly reduces the number of trainable parameters for downstream tasks while preserving the adaptability of the learned representations. The fine-tuning architecture and settings are similar to those in MLP probing, with only the transformer’s ${W_{q},W_{k},W_{v},W_{o}}$ parameters fine-tuned.
@@
Figure 2A provides a comprehensive visualization of performance across all benchmarked protein-related tasks. Distinct colors and shapes in the figure represent different evaluation strategies and models, respectively. The ESM2-150M model, being relatively smaller, serves as an indicator to assess the difficulty of these tasks. The performance distribution highlights the inherent relationships between the complexity of tasks and the advantages brought by the scale of the model. The distribution of performance across tasks underscores the relationship between task complexity and the benefits derived from the scale of the model. In more complex tasks, such as Contact Map prediction (Protein Structure category), Fluorescence (Protein Function), Metal Binding (Protein Interaction), and Stability (Protein Development), the larger models (xTrimoPGLM-100B and ESM2-15B) significantly outperform the smaller ESM2-150M. This disparity in performance highlights the necessity for more advanced models to effectively address complex tasks. Conversely, for simpler tasks, like Antibiotic Resistance (Protein Function category), the performance gap between the large and small models is notably smaller. This observed pattern suggests that larger models are more adept at capturing the intrinsic evolutionary information of protein sequences. Consequently, as models scale up, they exhibit marked improvements in performance, particularly for complex tasks. The application of LoRA consistently enhances overall task performance compared to the static Probing method, which limits the capacity of pre-trained models (Figure 2B). LoRA’s efficiency lies in its ability to refine and utilize key features without significantly increasing the number of trainable parameters (less than 1% increase with LoRA rank set to 8). The empirical findings also highlight a scaling trend in task performance with supervised fine-tuning, suggesting a strong correlation between model scale and performance. The following sections will delve deeper into the varying scaling behaviors observed across different types of tasks.
Scaling Behaviors of Downstream Tasks. We extend the investigation of scaling laws to downstream tasks in four distinct categories of protein-related tasks from three perspectives.
@@
Comparison Between ESM and xTrimoPGLM Family models. We observe the scaling behavior of downstream tasks among the ESM and xTrimoPGLM family models using both Linear Probing with MLP (Pb) and Fine-tuning with LoRA (Ft) techniques (Figure 2B). The x-axis measures the total computational cost in Petaflop Days (PF-Days), while the y-axis represents the performance of various task categories. The results indicate that, although the intensity of the scaling effect varies among task types, a common trend persists: an exponential increase in computational resources during pre-training correlates with linear improvements in downstream task performance. While the intensity of the scaling effect differs among task types, a common trend observed from pre-training extends to these downstream tasks: an exponential increase in computational resources during pre-training correlates with linear improvements in task performance. This extension of the scaling law to downstream tasks in protein modeling is a novel observation.
@@
Comparison Among Different Training States of xTrimoPGLM. To eliminate the differences in backbone architectures, pre-train datasets, etc, we further take a deep insight into the correlations between the downstream task performance with the ongoing pre-training process (measured by the MLM validation loss, Figure 2C). We observe that most tasks demonstrate positive correlations. Specifically, 44% show a gradual increase in performance, 28% exhibit a “Grokking” phenomenon, and 16% do not follow a strictly linear pattern. The ’Grokking’ phenomenon mirrors emergent abilities seen in large NLP models  (?). It occurs when models initially overfit the training data, especially when task datasets have limited overlap with pre-training data. As the model’s knowledge base expands, it begins to apply its understanding to out-of-distribution (OOD) scenarios, resulting in a sudden increase in test data performance. Thus the performance of the test data suddenly increased, denoted as the Grokking phenomenon. However, 12% of tasks indicate a potential negative impact from increased computational resources, suggesting an inverse scaling effect.
Comparison Among Different Scales of xTrimoPGLM Family Models. Additionally, we further analyze the scaling behavior by comparing xTrimoPGLM family models in the Supplementary Figure 7, which show a similar trend that most tasks exhibit a positive relationship between task performance and the scaling of training FLOPs and model size.
@@
Figure 3: Structure Prediction with xT-Fold. A. xT-Fold architecture leverages a Multi-Layer Perceptron (MLP) to convert PLM representations into inputs for the folding modules, which generate 3D coordinates and pLDDT confidence scores. B. TM-score benchmarks for structure prediction models. The bar chart shows the performance of single-sequence PLM-based models and MSA-based models on CAMEO and CASP15 datasets. C. Inference time comparison across models for varying sequence length intervals, showing xT-Fold, ESMFold, OmegaFold, AlphaFold, RoseTTAFold, and MSA search times in seconds.D. Scatter plots compare xT-Fold predictions (x-axis) to other models (y-axis), color-coded by perplexity (green for high, purple for low).